Find the input thermal noise voltage of a receiver with a bandwidth of 3.33 kHz, with an input resistance of 42 Ω, and a temperature of 29oC.
1
Expert's answer
2021-10-10T16:05:56-0400

Noise voltage are unwanted and disturbing voltage in an electronic system. It is given by:

"V^2 = 4kT \int^{f2}_{f1} Rdf"

where V = RMS input noise voltage

"f_2-f_1 = Band \; width = 3.33 \;kHz = 3.33 \times 10^3 \;Hz \\

k = Boltzmann \;constant = 1.38 \times 10^{-23} \;J/K \\

T = Temperature = 29+273 = 302 \;K

R = Resistance = 42 \; \Omega"

Putting all these values we get

"V = (4kTR(f_2 -f_1))^{0.5} \\

V = (4 \times 1.38 \times 10^{-23} \times 302 \times 42 \times 3.33 \times 10^3)^{0.5} \\

V = 482.86 \times 10^{-10} \;V"